22 Question (4 points) What is the percent composition of the antibiotic tetracycline, which has the molecular formula of C22Part 2 (1 point) % H in C22H24N208 Part 3 (1 point) % N in C22H24N208 Part 4 (1 point) % 0 in C22H24N208
Answer #1

Part 1 : 59.45 % C in C22H24N2O8

Part 2 : 5.44 % H in C22H24N2O8

Part 3 : 6.30 % N in C22H24N2O8

Part 4 : 28.80 % O in C22H24N2O8

Explanation

Part 1

molar mass C22H24N2O8 = 444.4346 g/mol

mass C in C22H24N2O8 = 12 * (molar mass C)

mass C in C22H24N2O8 = 12 * (12.0107 g/mol)

mass C in C22H24N2O8 = 144.1284 g/mol

mass % C = (mass C in C22H24N2O8 / molar mass C22H24N2O8) * 100

mass % C = (144.1284 g/mol / 444.4346 g/mol) * 100

mass % C = (0.5945) * 100

mass % C = 59.45 %

New Microsoft Word Document.docx- Microsoft Word (Product Activation Failed) X Page Layout File Home Insert References Mailin